OpenStudy (anonymous):

How many triangles ABC satisfy the following conditions? A=50°, b = 19 ft, a = 9 ft

Directrix (directrix):

Given two sides and a non-included angle is known as the Ambiguous Case. There may be 0, 1, or 2 triangles formed with such data. Using the Law of Sines as a beginning point: 9/ sin(50) = 19/sin B 11.749 = 19/sinB sin B = 19/11.749 sin B = 1.617 which exceeds 1. The largest value of sine is 1. Therefore, no such triangle exists with this data.

OpenStudy (anonymous):

So the answer is 0 or impossible ?

Directrix (directrix):

It is impossible to have a triangle with those measurements. So, there are zero such triangles. Use whatever language you use in class.

OpenStudy (anonymous):

Okay thanks Directrix :D